# Phonon, Infrared and Raman Spectra of LiGa5O8 from Density Functional Perturbation Theory

### v1

@@ I INTRODUCTION
LiGa5O8 was recently reported to be a p-type ultra-wide-band-gap (UWBG) semiconductor with a gap of 5.34 eV and cubic spinel structure [1]. This is an unusual finding as native p-type doping in UWBG oxides is rare and p-type doping is often hindered by small polaron formation. The experimental gap [1] extracted from a Tauc plot is slightly smaller than the calculated gap of 5.72 eV [2] using the quasiparticle self-consistent (QS) method, which is a many-body perturbation theory (MBPT) method in which the self-energy is calculated from the one particle Green’s function and the screened Coulomb interaction including electron-hole interactions in the screening. The optical band gap obtained from the Bethe Salpeter Equation (BSE) in this same work [2] indicates a large exciton binding energy with lowest direct exciton gap of 5.5 eV and also showed the existence of a 0.1 eV lower indirect gap. The band gap is thus fairly well established. However, the origin of the p-type doping is not. Native defect calculations [3] showed all plausible acceptors to have high binding energies exceeding 0.75 eV. Furthermore, compensating shallow donors like GaLi antisites or deep donor oxygen vacancies were found to pin the equilibrium Fermi level deep in the gap thus predicting insulating behavior. Similar findings were found with evidence of small polaron formation for native defects and some candidate acceptor impurities by Lyons [4]. Insulating behavior has meanwhile been found upon annealing in O2 and the stoichiometry conditions under which p-type behavior occurs have been more clearly established [5] to be somewhat O-poor, Li-poor and Ga-rich. In spite of the lack of understanding of the origin of p-type doping rectification behavior at pn-heterojunçtions between n-type Ga2O3 and p-type LiGa5O8 has been reported [6]. LiGa5O8 has previously received some interest as a host for phosphorescent dopants like Cr [7, 8, 9]. The spinel structure of this compound was reported by Joubert et al.[10]. While the present paper does not address directly the question of p-type doping, the renewed interest in this material motivates a broader study of its fundamental properties. Here we present a first-principles study of the phonons with the aim to assist on-going characterization efforts on this new material. First, the phonon band structure calculations are important to verify the mechanical stability of the material in this structure. Second, Raman spectra are often used to characterize the crystalline quality of a material. Third, electron-phonon coupling is responsible for finite temperature gap renormalization and for limitations to the mobility. We here focus on the phonons at the Brillouin zone center and present a symmetry analysis of the vibrational modes of the 56 atom unit cell and report the calculated Infrared and Raman spectra as well as related Born effective charges and static and high frequency dielectric constants.
@@ III.4 Phonon dispersions and density of states
It shows the usual quadratic onset in the region of acoustic modes. The decomposition in atomic components shows that the highest modes have primarily O character. Even though Li atoms have smaller mass, their main contributions appear between 300-400 and near 500 cm-1. This indicates that the Li-O bonds have smaller force constants than some of the Ga-O ones. This is consistent with the bond-lengths in Table 1. The lower energy range up to 250 cm-1 shows primarily Ga contribution as expected because Ga is the heaviest atom in the system.
@@ III.6 IR Spectra and Dielectric Properties
The poles of the dielectric function, which correspond to peaks in its imaginary part correspond to the TO modes, while the zeros of the real part correspond to the LO modes. The latter also show up as peaks of the loss function . From these, one can extract the absorption coefficient with the complex index of refraction and the normal incidence reflectance Using a fixed broadening parameter of 5 cm-1 for all modes, we obtain the dielectric function and related properties, displayed in fig- 5. Because some modes, notably modes 8(61-63), 11 (86-88), 16(117-119), have very small oscillator strength (and small TO-LO) splitting, they are not clearly visible and hence we see only 18 instead of the predicted 21 peaks. Note that the strongest mode is mode which corresponds to modes 69-71 when ordering all modes irrespective of symmetry according to energy as given in Table 5. As an example, we show the eigendisplacement pattern for mode in fig. 6. We can see that most of the cations have an upward motion while the O atoms have a downward motion. Thus, there is a significant dipole character to the overall motions consistent with a large IR active oscillator strength. There are both Li-O and Ga-O bond stretches involved in this mode and they are in phase for equivalent atoms.
@@ III.7 Raman spectra
In fig.8 we show the Raman modes of and symmetry. Note that the - modes have significanty smaller intensity. Both of these are active in a scattering geometry with parallel incident and scattered light polarizations. We show the spectra for modes separately in fig.8 because these are active for perpendicular incident and

### v2

@@ I INTRODUCTION
LiGa5O8 was recently reported to be a p-type ultra-wide-band-gap (UWBG) semiconductor with a gap of 5.34 eV and cubic spinel structure [1]. This is an unusual finding as native p-type doping in UWBG oxides is rare and p-type doping is often hindered by small polaron formation. The experimental gap [1] extracted from a Tauc plot is slightly smaller than the calculated gap of 5.72 eV [2] using the quasiparticle self-consistent (QS) method, which is a many-body perturbation theory (MBPT) method in which the self-energy is calculated from the one particle Green’s function and the screened Coulomb interaction including electron-hole interactions in the screening. The optical band gap obtained from the Bethe Salpeter Equation (BSE) in this same work [2] indicates a large exciton binding energy with lowest direct exciton gap of 5.5 eV and also showed the existence of a 0.1 eV lower indirect gap. The band gap is thus fairly well established. However, the origin of the p-type doping is not. Native defect calculations [3] showed all plausible acceptors to have high binding energies exceeding 0.75 eV. Furthermore, compensating shallow donors like GaLi antisites or deep donor oxygen vacancies were found to pin the equilibrium Fermi level deep in the gap thus predicting insulating behavior. Similar findings were found with evidence of small polaron formation for native defects and some candidate acceptor impurities by Lyons [4]. Insulating behavior has meanwhile been found upon annealing in O2 and the stoichiometry conditions under which p-type behavior occurs have been more clearly established [5] to be somewhat O-poor, Li-poor and Ga-rich. In spite of the lack of understanding of the origin of p-type doping rectification behavior at pn-heterojunçtions between n-type Ga2O3 and p-type LiGa5O8 has been reported [6]. LiGa5O8 has previously received some interest as a host for phosphorescent dopants like Cr [7, 8, 9]. The spinel structure of this compound was reported by Joubert et al.[10]. While the present paper does not address directly the question of p-type doping, the renewed interest in this material motivates a broader study of its fundamental properties. Here we present a first-principles study of the phonons with the aim to assist on-going characterization efforts on this new material. First, the phonon band structure calculations are important to verify the mechanical stability of the material in this structure. Second, Raman spectra are often used to characterize the crystalline quality of a material. Third, electron-phonon coupling is responsible for finite temperature gap renormalization and for limitations to the mobility. We here focus on the phonons at the Brillouin zone center and present a symmetry analysis of the vibrational modes of the 56 atom unit cell and report the calculated Infrared and Raman spectra as well as related Born effective charges and static and high frequency dielectric constants. In prior work, we studied the phonons and related properties in LiGaO2, [11] another compound in the Li-Ga-O system but that compound has a simpler 16 atom wurtzite based structure. Our results on that material obtained with a similar approach as here were in good agreement with experimental data.
@@ III.3 Electronic band structure
In Fig.3 we show a zoom in of the band structure near the valence band maximum (VBM). We can see that the VBM occurs at and the band is doubly degenerate along and on the Brillouin zone surface but splits along into a heavy hole and light hole band. These two bands near can be described by an effective Hamiltonian of the form
with the deviation from point , the unit matrix and the Pauli matrix. The inverse mass parameters are related to , which are the components of the effective mass tensor of the degenerate band in and direction, while give the inverse heavy and light hole masses along . The conduction band minimum (CBM) at has an isotropic mass tensor because of the cubic symmetry. These masses can be determined simply by fitting a behavior to the dispersion near the band extrema and lead to the values given in Table 3. We here give the masses both extracted from the current PBEsol band structure results and from the all-electron PBESol and QS results from [2]. Note that the heavy massses suffer from larger uncertainty due to the small changes of the band over the region where parabolic dispersion holds and are sensitive to the range of k used in the fit. However, within each set of results, the relation between the effective masses along and predicted by the effective Hamiltonian hold fairly well. PBEsol is seen to overestimate the masses relative to the more accurate QS. For the conduction band mass the density functional perturbation theory approach as implemented in Janssen et. al.[23] gives the same result as the direct fit and good agreement between the different approaches is obtained. For the VBM the double degeneracy complicates the calculation of the effective mass tensor. For in the -plane, we find a high effective mass along the -direction and for the top band along the [110] direction but there are 12 equivalent points at , , and hence for every there is some direction with heavy mass. These heavy masses found here are similar to those in Ga2O3[24, 25] and indicate that hole mobilities may be low and self-trapped polarons are likely to occur.
@@ III.4 Phonon dispersions and density of states
It shows the usual quadratic onset in the region of acoustic modes. The decomposition in atomic components shows that the highest modes have primarily O character. Even though Li atoms have smaller mass, their main contributions appear between 300-400 and near 500 cm-1. This indicates that the Li-O bonds have smaller force constants than some of the Ga-O ones. This is consistent with the bond-lengths in Table 1. The lower energy range up to 250 cm-1 shows primarily Ga contribution as expected because Ga is the heaviest atom in the system. A phonon DOS like spectrum might be expected in Raman spectra of samples with significant disorder breaking the crystal momentum conservation, while highly crystalline samples would show only zone centrum modes. Furthermore a reduction of phonon DOS in the regions 300-400 cm-1 and 500 cm-1 in observed spectra might indicate the presence of Li vacancies while an enhancement in the low energy region might indicate Ga rich stoichiometry or a reduction in this low energy region, Ga vacancies. A reduction of the highest region of the spectra might be indicative of oxygen vacancies. Thus, the Raman spectra might provide indirect information on the defects present in the material.
@@ III.6 IR Spectra and Dielectric Properties
The poles of the dielectric function, which correspond to peaks in its imaginary part correspond to the TO modes, while the zeros of the real part correspond to the LO modes. The latter also show up as peaks of the loss function . From these, one can extract the absorption coefficient with the complex index of refraction and the normal incidence reflectance Using a fixed broadening parameter of 5 cm-1 for all modes, we obtain the dielectric function and related properties, displayed in Fig. 6. This is a typical value used to obtain a smooth spectrum. The actual phonon linewidth may vary from mode to mode and has contributions from electron-phonon and three- and four-phonon scattering anharmonic terms, and are here not calculated [26, 27, 28]. When experimental spectra become available, one could adjust the linewidth to obtain a better fit but in the absence of such information, the single value used here is sufficient to provide a first approximation to the spectral shape. Because some modes, notably modes 8(61-63), 11 (86-88), 16(117-119), have very small oscillator strength (and small TO-LO) splitting, they are not clearly visible and hence we see only 18 instead of the predicted 21 peaks. Note that the strongest mode is mode which corresponds to modes 69-71 when ordering all modes irrespective of symmetry according to energy as given in Table 6. As an example, we show the eigendisplacement pattern for mode in Fig. 7. We can see that most of the cations have an upward motion while the O atoms have a downward motion. Thus, there is a significant dipole character to the overall motions consistent with a large IR active oscillator strength. There are both Li-O and Ga-O bond stretches involved in this mode and they are in phase for equivalent atoms.
@@ III.7 Raman spectra
The Raman simulated spectra are obtained by broadening the modes into a Gaussian peak with the same uniform linewidth of 5 cm-1 as used for the IR spectra. Again, this is just a typical value and may require adjustment for a better fit to experiment. In Fig.10 we show the Raman modes of and symmetry. Note that the - modes have significantly smaller intensity. Both of these are active in a scattering geometry with parallel incident and scattered light polarizations. We show the spectra for modes separately in Fig.10 because these are active for perpendicular incident and
at 715.1 cm-1
@@ Acknowledgements.
- Boonchun and Lambrecht [2010] A. Boonchun and W. R. L. Lambrecht, First-principles study of the elasticity, piezoelectricity, and vibrational modes in compared with ZnO and GaN, Phys. Rev. B 81, 235214 (2010).
@@
- Laflamme Janssen et al. [2016] J. Laflamme Janssen, Y. Gillet, S. Poncé, A. Martin, M. Torrent, and X. Gonze, Precise effective masses from density functional perturbation theory, Phys. Rev. B 93, 205147 (2016).
- Poncé and Giustino [2020] S. Poncé and F. Giustino, Structural, electronic, elastic, power, and transport properties of from first principles, Phys. Rev. Res. 2, 033102 (2020).
- Mock et al. [2017] A. Mock, R. Korlacki, C. Briley, V. Darakchieva, B. Monemar, Y. Kumagai, K. Goto, M. Higashiwaki, and M. Schubert, Band-to-band transitions, selection rules, effective mass, and excitonic contributions in monoclinic , Phys. Rev. B 96, 245205 (2017).
- Han et al. [2022a] Z. Han, X. Yang, S. E. Sullivan, T. Feng, L. Shi, W. Li, and X. Ruan, Raman Linewidth Contributions from Four-Phonon and Electron-Phonon Interactions in Graphene, Phys. Rev. Lett. 128, 045901 (2022a).
- Li et al. [2014] W. Li, J. Carrete, N. A. Katcho, and N. Mingo, Shengbte: A solver of the boltzmann transport equation for phonons, Computer Physics Communications 185, 1747 (2014).
- Han et al. [2022b] Z. Han, X. Yang, W. Li, T. Feng, and X. Ruan, FourPhonon: An extension module to ShengBTE for computing four-phonon scattering rates and thermal conductivity, Computer Physics Communications 270, 108179 (2022b).